# Optimal Parallel $\lambda$-Reduction

### v1

@@ 4.  From $\Delta$-nets back to $\lambda$-terms
Meging unpaired replicators always reduces the total number of interactions and the total number of agents, improving space and time efficiency. The optimal partial ordering among replicator commutations is one where active pairs with unpaired replicators are reduced last (specially active pairs of two unpaired replicators). This gives a chance for replicators of unknown status to accumulate behind unpaired replicators and potentially be determined to be unpaired, which would allow them to be merged together. Unpaired replicators can be merged lazily, as part of the fan replication and replicator commutation core interaction rules.
@@
    Before applying every commutation with an unpaired replicator:
    • $\bullet$
      it should be merged with all unpaired replicators which are connected to its auxiliary ports (unpaired replicator merging) [I,K]*
    • $\bullet$
      all auxiliary ports connected to erasers should be removed (partial unpaired replicator decay) [K]*

### v2

@@ 4.  From $\Delta$-nets back to $\lambda$-terms
Before applying a rule that involves an unpaired replicator (other than replicator erasure), it is always best to merge the unpaired replicator with consecutive unpaired replicators and decay it if possible. Relative to a baseline reduction order in which replicators are only merged when no other reductions are possible, this can only reduce the total number of reductions and the total number of agents, improving space and time efficiency. Unpaired replicators can be merged lazily, immediately before the fan replication, replicator replication, and aux fan replication rules. When multiple rules can be applied, it is advantageous to apply rules involving unpaired replicators last (specially those with two unpaired replicators) as this allows replicators of unknown status to accumulate behind unpaired replicators and potentially be determined to be unpaired and then merged.
@@
Before applying fan replication, replicator replication, or aux fan replication, each involved unpaired replicator should:
  • $\bullet$
    be merged with all unpaired replicators which are connected to its auxiliary ports (unpaired replicator merging) [I,K]*
  • $\bullet$
    have all auxiliary ports connected to erasers removed (partial unpaired replicator decay) [K]*